# The GPT-WritingPrompts Dataset: A Comparative Analysis of Character Portrayal in Short Stories

### v1

@@ 4.4 Converting Attributes to Scores
The final, crucial component of the pipeline is to convert our extracted list of protagonist attributes to numerical scores along each of our selected dimensions of portrayal. We condense the methods proposed in past literature for various dimensions into a generalized set of algorithms, enabling a comparative analysis of the consistency of scores using these different methods.
Average lexical score (lex-avg)
: We start again with a simple baseline: the average of the lexical scores, for attribute tokens that are present in the lexicons for each of our dimensions.
Average embedding similarity (emb-sim)
: Here, attribute tokens are not in the lexicon for a dimension are assigned scores based on the average cosine similarity with the lexicon terms. More concretely, given the dimension lexicon $L$, and the protagonist attribute $t$, we calculate the dimensional intensity of $t$ as:
@@
Where $e(\cdot)$ is the pretrained word2vec embedding Mikolov et al. (2013). This is the process followed in Lucy and Bamman (2021) and Huang et al. (2021) for the Appearance and Intellect dimensions. We adapt it for the VAD dimensions by considering cosine similarity to words that are above the 75th percentile of scores in the respective lexicons, and for Power by considering only those words that are associated with high power. The average of the scores of all attribute tokens represents the score for a story.
Axis projection (axis-emb)
: Here, we construct a vector axis, or semantic axis, that encodes the dimensional property as a direction in the embedding space. Given a set of terms $L_{l}$ and $L_{h}$ from $L$ that represent the extreme ends of the dimension – say, a list of low valence and high valence terms – we find the vector axis for the dimension by computing the following:
@@
For a given attribute token $t$, we then find its score $S(t,L)$ by projecting it’s embedding onto this vector axis, i.e, computing the cosine similarity between the attribute embedding and the axis. The average of the scores of all attribute tokens represents the score for a story. Note that here, we need a bipolar dimension, with terms that represent the lower end and higher end of the dimension. This approach has been used in prior work to compute scores along the Power dimension, and we adapt it for the VAD dimensions as well [^4] High and low scoring terms are defined by those scoring above the 75th and below the 25th percentile of the lexicon.; however, it cannot be applied to the Appearance and Intellect dimensions.
@@ 4.4.1 Evaluation
We first conduct an evaluation of the methods for estimating dimension scores given a list of attribute tokens, using the annotated lexicons for the three VAD dimensions.
We divide the lexicon for each dimension into a train and test split, and estimate the scores for terms in the test split using the emb-sim and axis-emb methods. We evaluate each method by measuring the Spearman correlation between the assigned scores from each method and the gold scores from the VAD lexicon for terms in the test split. Results for each of the VAD dimensions are in Table 2.
We see that the axis-emb scoring method is substantially better at estimating the intensity scores of non-lexicon terms compared to the emb-sim method, for all three dimensions. We therefore perform all subsequent analyses with axis-emb scores for the bi-polar dimensions of valence, arousal, dominance, and power. For appearance and intellect, we are only estimating a uni-polar score of the strength of association with appearance and intellect related terms, and therefore use the emb-sim method.
Figure 1: The proportion of stories that fall under the five inferred point-of-view (PoV) categories for stories written by humans and generated by GPT-3.5, grouped by the inferred PoV specified by the prompt.
@@ 5.1 Distribution of Protagonist Groups
We first analyze the distribution of stories for each PoV category, for the human-written and GPT-generated subsets of our dataset.
We label each input prompt as belonging to a PoV category by the pre-dominant pronominal tokens used in them, analogous to stories.
Figure 1 shows the proportion of generated stories that fall into each PoV category, for each of the PoVs categories of prompts. We see that GPT-generated stories largely tend to be in the third-person, irrespective of the PoV specified by the prompt. For prompts in the third-person, a majority of the generated stories are consistent with the protagonist gender specified in the prompt (TP-M and TP-F), whereas human-written stories for the same prompts show more diversity in their chosen protagonist genders. With first and second-person prompts (FP and SP), the model favors generating third-person stories with male protagonists over female protagonist. This bias also holds for human-written stories, though the majority of stories by the latter are in first-person.
Figure 2: Distribution of z-scored COMET (shaded dark) and spaCy (lightly shaded) attribute scores along each dimension for the different PoV groups, and human and gpt-3.5 generated stories. Word-level scores are calculated with the axis-emb method for VAD and Power, and emb-sim for appearance and intellect.
@@ 5.2 Variation with Point-of-View
We now compare the scores of protagonist attributes extracted using spaCy and COMeT for each PoV group, and between human-written and GPT-generated stories. We use the axis-emb method to score attribute terms for the bi-polar VAD dimensions, and the emb-sim method for the uni-polar appearance and intellect dimensions. We normalize all the scores along each dimension (across all human and machine-generated stories) via z-scoring.
@@
Results: Figure 2 plots the mean score for each of the PoV and writer groups (standard deviation is indicated with error bars) (we also report the mean and standard deviation of COMET-attribute scores for all groups in Table 3).
@@
Across all PoV categories, GPT-generated stories portray their protagonists with a higher valence, lower arousal, and higher dominance (and power) when compared to human-written stories (all differences are statistically significant at p$<$0.00). Generated stories use fewer words associated with both appearance and intellect for their protagonists when compared to human writers. From Table 3, we also find that human-written stories have a higher standard deviation in scores across all dimensions and groups when compared to model generations, indicating that the former group demonstrates a higher variability in language production when compared to neural generators Giulianelli et al. (2023); however, a rigorous evaluation of this hypothesis requires a much larger sample of stories per prompt than in our dataset currently, and we leave it for future work to explore.
@@
Comparing the portrayal of male and female protagonists, we find that TP-F stories have higher valence, and lower arousal and dominance scores associated with them for both human and machine-generated stories. TP-F stories also use more words associated with appearance, and slightly fewer words associate with intellect, when compared to TP-M stories. These differences are statistically significant at p$<$0.001.
@@
Comparison with spaCy attributes: Figure 2 also plots the mean portrayal for each PoV group when protagonist attributes are extracted using linguistic dependencies with spaCy in lighter shades of markers. We observe that trends in differences in portrayal largely follow those demonstrated by COMET attributes: TP-F stories have higher scores for valence, lower scores for dominance and arousal, and use more words associated with appearance. However, we note that the relative levels of intensity for the different PoV groups do not always align with those obtained with COMET attributes, particularly for the appearance and intellect dimensions. Stories that fall under the "Other" category, for example, have a higher mean association score with intellect-associated words when compared to FP stories when compared using COMET attributes, but this trend is notably reversed for GPT-generated stories when analyzed with spaCy attributes. We hypothesize that this is indicative of syntactic differences in how character attributes are expressed in different types of stories, and consider it an interesting avenue of exploration for future work.
@@ 5.3 Variation with Prompts
The previous sections considered differences in the mean and standard deviation of the scores across all the stories under a PoV label. We now consider the differences in scores when conditioned by the input prompt, i.e, how do the two text generation processes (human and gpt-3.5) vary in their outputs when prompted with the same input context?
We study this by computing instance-level differences in the portrayal of the stories generated by each writer group, where each prompt is considered to be an instance. The prompt-score is estimated as the mean of the portrayal scores (using COMET attributes) of the stories associated with that prompt. In order to keep sample sizes consistent between the two writer groups, we compute the prompt-score for the human subset by sampling two stories at a time, and further averaging this across five sampling runs. We then compute the prompt-wise differences in scores between human writings and GPT generations. In order to contextualize these differences, we use a human control group: the human-written stories for each prompt are split into two mutually-exclusive subsets, and prompt-wise differences between a human split and GPT generations, and between the two human splits.
@@
Results: In Figure 3(a), we plot the distribution of the instance-level score differences between human-GPT stories. The differences follow a normal distribution centered close to 0 for all dimensions, and deviations are equally distributed on both sides of the mean; for valence, dominance, and power, the median difference falls below 0, inline with the results in Figure 2 that show GPT-generated stories as having a lower mean valence and dominance across all PoV groups (similar results hold for the arousal, appearance, and intellect dimensions).
Taking advantage of the normality of distributions of these differences, we summarize each distribution by its mean value. In Figure 3(b), we plot the mean of the differences between human and gpt-3.5 generations, along with the mean of the differences between the two human control splits. Human-written stories are significantly closer to one another in how they portray their protagonists in response to a prompt when compared to GPT-generations. These portrayal scores therefore provide a strong signal in distinguishing GPT-generated stories from those written by humans. This adds to, and is inline with, a growing body of work in NLP that demonstrates that neural text generators are still far from imitating human models of language production Giulianelli et al. (2023); Chakrabarty et al. (2023).
@@ 6 Conclusion
Neural language models are increasingly being used in text generation tasks that span diverse use-cases, ranging from controlled generation tasks like summarization, question answering, and stylized translations to open-ended creative generation tasks. In this work, we studied the properties of one such generative model, gpt-3.5, in the context of a controlled creative generation task: writing a short story in response to a short prompt. We expanded the WritingPrompts dataset with gpt-3.5 generations to create the GPTWritingPrompts dataset, and used this to compare the GPT generative model with human models of language production in response to the same prompts.
Our analysis shows that LLM-generated stories tend to be more positive, less active (lower arousal), and more in-control (higher dominance) when compared to human-written stories, and also demonstrate less variation in these scores across their generations. We also examined differences in the portrayal of protagonists in these stories when grouped by the narrative style (point-of-view) and inferred gender of the main character, and found that both machine and human generations display significant and consistent biases in how female protagonists are portrayed (higher valence, lower arousal, lower dominance, higher associations with appearance, and lower associations with intellect) when compared to male protagonists. These results demonstrate how neural generative models can both differ from, and imitate, human models of language production. Our dataset of paired generations is also a useful resource for researchers working on studying these differences both from the perspective of generative models of text and in relation to biases in open-ended generation.
7 Limitations
@@ Appendix A Additional Tables
In Table 4, we list a human-written story and a GPT-generated response to a sample of three prompts from the WritingPrompts dataset. The stories are off at 100 tokens to be compact.

### v2

@@ 4.4 Converting Attributes to Scores
The final, crucial component of the pipeline is to convert our extracted list of protagonist attributes into numerical scores along each of our selected lexical dimensions. We condense the methods proposed in past literature into a generalized set of algorithms, enabling a comparative analysis of the consistency of results using different methods.
Average lexical score (lex-avg): We start with a simple baseline: the average of the lexical scores for attribute tokens that are directly present in the lexicons for each of our dimensions (assigning a score of 1 to the token smart, which is directly in the lexicon for Intellect, for example).
Average embedding similarity (emb-sim): Here, attribute tokens not in the lexicon for a dimension are assigned scores based on the average cosine similarity with the lexicon terms. More concretely, given the dimension lexicon $L$, and the protagonist attribute $t$, we calculate the dimensional intensity of $t$ as:
@@
Where $e(\cdot)$ is the pretrained word2vec embedding Mikolov et al. (2013). This is the process followed in Lucy and Bamman (2021) and Huang et al. (2021) for the Appearance and Intellect dimensions. We adapt it for the VAD dimensions by considering cosine similarity to words $L$ that are above the 75th percentile of scores in the respective lexicons, and for Power by considering only those words that are associated with high power. The average of the scores of all attribute tokens represents the score for a story.
Axis projection (axis-emb): Here, we construct a vector axis, or semantic axis, that encodes the dimensional property as a direction in the embedding space. Given a set of terms $L_{l}$ and $L_{h}$ from $L$ that represent the extreme ends of the dimension – say, a list of low valence and high valence terms – we first find the vector axis for the dimension:
@@
For a given attribute token $t$, we then find its score $S(t,L)$ by projecting it’s embedding onto this vector axis, i.e, computing the cosine similarity between the attribute embedding and the axis. The average of the scores of all attribute tokens represents the score for a story. Note that here, we need a bipolar dimension, with terms that represent the lower end and higher end of the dimension. This approach has been used in prior work to compute scores along the Power dimension, and we adapt it for the VAD dimensions  [^5] High and low scoring terms are defined by those scoring above the 75th and below the 25th percentile of the lexicon..
@@ 4.4.1 Evaluation
We divide the lexicon for each VAD dimension into a train and test split, and estimate the scores for terms in the test split using the emb-sim and axis-emb methods. We evaluate each method by measuring the Spearman correlation between the assigned scores from each method and the gold scores from the VAD lexicon for terms in the test split. Results are reported in Table 2.
We see that the axis-emb scoring method is substantially better at estimating the intensity scores of non-lexicon terms compared to the emb-sim method, for all three dimensions. We therefore perform all subsequent analyses with axis-emb scores for the bipolar dimensions of valence, arousal, dominance, and power. For appearance and intellect, we are only estimating a unipolar score of the strength of association with appearance and intellect related terms, and therefore use emb-sim.
Figure 1: The proportion of stories that fall under the five inferred point-of-view (PoV) categories for stories written by humans and generated by GPT-3.5, grouped by the inferred PoV of the prompt.
@@ 5.1 Distribution of Protagonist Groups
We first analyze the distribution of stories for each PoV category, for the human-written and GPT-generated subsets of our dataset. We label each input prompt as belonging to a PoV category by the pre-dominant pronominal tokens used in them, analogous to stories.
Results: Figure 1 shows the proportion of generated stories that fall into each PoV category, for each of the PoVs categories of prompts. We see that GPT-generated stories largely tend to be in the third-person, irrespective of the PoV specified by the prompt. For prompts in the third-person, a majority of the generated stories are consistent with the protagonist gender specified in the prompt (TP-M and TP-F), whereas human-written stories for the same prompts show more diversity in their chosen protagonist genders. With first and second-person prompts (FP and SP), the model favors generating third-person stories with male protagonists over female protagonists. This bias also holds for human-written stories, though the majority of stories by the latter are in first-person.
Figure 2: Distribution of z-scored COMET (shaded dark) and spaCy (lightly shaded) attribute scores along each dimension for the different PoV groups, and human and GPT-3.5 generated stories.
@@ 5.2 Variation with Point-of-View
We now compare the scores of protagonist attributes extracted using spaCy and COMeT for each PoV group, and between human-written and GPT-generated stories. We normalize all the scores along each dimension (across all human and machine-generated stories) via z-scoring.
@@
Results: Figure 2 plots the mean score for each of the PoV and writer groups (standard deviation is indicated with error bars). We also report the mean and standard deviation of COMeT-attribute scores for all groups in Table 3.
@@
Across all PoV categories, GPT-generated stories portray their protagonists with a higher valence, lower arousal, and higher dominance (and power) when compared to human-written stories (all differences are statistically significant at p$<$0.001). Generated stories use fewer words associated with both appearance and intellect for their protagonists when compared to human writers. From Table 3, we also find that human-written stories have a higher standard deviation in scores across all dimensions and groups when compared to model generations, indicating that the former group demonstrates a higher variability in language production when compared to neural generators Giulianelli et al. (2023); Breithaupt et al. (2024); however, a rigorous evaluation of this hypothesis requires a much larger sample of stories per prompt than in our dataset currently, and we leave it for future work to explore.
@@
Comparing the portrayal of male and female protagonists, we find that TP-F stories have higher valence, and lower arousal and dominance scores associated with them for both human and machine-generated stories. TP-F stories also use more words associated with appearance, and slightly fewer words associate with intellect, when compared to TP-M stories, in both sets of stories. These differences are statistically significant at p$<$0.001.
@@
Comparison with spaCy attributes: Figure 2 also shows the mean portrayal for each PoV group when protagonist attributes are extracted using linguistic dependencies with spaCy. We observe that trends in differences in portrayal largely follow those demonstrated by COMET attributes: TP-F stories have higher scores for valence, lower scores for dominance and arousal, and use more words associated with appearance. However, we note that the relative levels of intensity for the different PoV groups do not always align with those obtained with COMET attributes, particularly for the appearance and intellect dimensions. Stories that fall under the "Other" category, for example, have a higher mean association score with intellect-associated words when compared to FP stories with COMET attributes, but this trend is notably reversed for GPT-generated stories when analyzed with spaCy attributes. We hypothesize that this is indicative of syntactic differences in how character attributes are expressed in different types of stories, and consider it an interesting avenue of exploration for future work.
@@ 5.3 Variation with Prompts
The previous sections examined differences in the mean and standard deviation of portrayal scores across all the stories under a PoV label. We now consider the differences in scores when conditioned by the input prompt, i.e, how do the two text generation processes (human and GPT-3.5) vary in their outputs when prompted with the same input context?
We measure this by computing instance-level differences in character portrayals scores for the stories generated by each writer group, where each prompt is considered to be an instance. The prompt-score is estimated as the mean of the portrayal scores (using COMET attributes) of the stories associated with that prompt. In order to keep sample sizes consistent between the two writer groups, we compute the prompt-score for the human subset by sampling two stories at a time, and further averaging this across five sampling runs. For each prompt, we then compute the difference between the average attribute scores of human stories and GPT generations. To contextualize the strength of these differences, we create a human control split: for each prompt, the human-written stories are split into two mutually-exclusive subsets, and instance-level score differences are computed between the two human story subsets (control), and between the human and GPT story subsets.
@@
Results: Figure 3(a) shows the distribution of the instance-level score differences between human-GPT stories. The differences follow a normal distribution centered close to 0 for all dimensions, and deviations are equally distributed on both sides of the mean; for valence, dominance, and power, the median difference falls below 0, inline with the results in Figure 2 that show GPT-generated stories as having a lower mean valence and dominance across all PoV groups (similar results hold for the arousal, appearance, and intellect dimensions).
Given the distributional normality of these differences, we summarize each distribution by its mean. In Figure 3(b), we plot the mean of the differences between human and gpt-3.5 generations, along with the mean of the differences between the two human control splits. Human-written stories are significantly closer to one another in how they portray their protagonists in response to a prompt when compared to GPT-generations. These portrayal scores therefore provide a strong signal in distinguishing GPT-generated stories from those written by humans. This adds to, and is inline with, a growing body of work in NLP that demonstrates that neural text generators are still far from imitating human models of language production Giulianelli et al. (2023); Chakrabarty et al. (2023); Breithaupt et al. (2024).
@@ 6 Conclusion
Neural language models are increasingly being used for a diverse range of text generation tasks, from summarization, question answering, and stylized translation, to open-ended creative generation. In this work, we expanded the Reddit WritingPrompts dataset with paired gpt-3.5-turbo generations to create the GPT-WritingPrompts dataset, and use this to compare generative and human models of language production in response to the same prompts.
We showed that LLM-generated stories tend to be more positive, less active, and more in-control when compared to human-written stories, and also demonstrate less variation in these scores across their generations. We also examined differences in the portrayal of protagonists in these stories when grouped by the narrative style (point-of-view) and inferred gender of the main character, and found that both machine and human generations display significant and consistent biases in how female protagonists are portrayed (higher valence, lower arousal, lower dominance, higher associations with appearance, and lower associations with intellect) when compared to male protagonists. These results demonstrate how neural generative models can both differ from, and imitate, human models of language production. Our dataset is also a useful resource for researchers working on studying generative models of text in relation to biases in open-ended generation.
Limitations
@@
Acknowledgements
The first author is supported by funding and resources provided by the Vector Institute, Toronto, and NAVER Labs.
@@ Appendix A Additional Tables
In Table 4, we list a human-written story and a GPT-generated response to a sample of three prompts from the WritingPrompts dataset. The stories are cut off at 100 tokens to be compact.